### v1

# A Behaviour and Disease Model of Testing and Isolation

Abstract
There has been interest in the interactions between infectious disease dynamics and behaviour for most of the history of mathematical epidemiology. This has included consideration of which mathematical models best capture each phenomenon, as well as their interaction, but typically in a manner that is agnostic to the exact behaviour in question. Here, we investigate interacting behaviour and disease dynamics specifically related to behaviours around testing and isolation. This epidemiological-behavioural interaction is of particular interest as, prospectively, it is well-placed to be informed by real-world data temporally monitoring test results and compliance with testing policy. To carry out our investigation we extend an existing “behaviour and disease” (BaD) model by incorporating the dynamics of symptomatic testing and isolation. We provide a dynamical systems analysis of the ordinary differential equations that define this model, providing theoretical results on its behaviour early in a new outbreak (particularly its basic reproduction number) and endemicity of the system (its steady states and associated stability criteria). We then supplement these findings with a numerical analysis to inform how temporal and cumulative outbreak metrics depend on the model parameter values for epidemic and endemic regimes. As the presented interdisciplinary modelling approach can accommodate further extensions (including, but not limited to, adding testing capacity, decay in behavioural effects and multiple pathogen variants), we hope that our work will encourage further modelling studies integrating specific measured behaviours and disease dynamics that may reduce the health and economic impacts of future epidemics.
Keywords Behavioural contagion  $\cdot$ Test, Trace and Isolate (TTI)  $\cdot$ Respiratory infection
1 Introduction
1.1 Background
From early in the history of infectious disease modelling, it was clear that the broad mathematical approaches used to model pathogen transmission could also be used to model social transmission such as the spread of rumours or behaviours, and that there were important specific differences between social and viral spreading, theoretically and empirically [1, 2]. As data availability and quality from outbreaks improved, empirical evidence for the impact of behavioural response on infection spread increased, particularly during the 2004 SARS epidemic [3]. A particularly influential 2009 paper by Funk et al. [4] (with over 800 citations at the time of writing) explicitly coupled mechanistic models of infection and behavioural response to infection. Even in 2010, a systematic review of approaches to this problem had 85 citations in its bibliography [5].
In the last 15 years there has been significant work on the interaction of behaviour and epidemics including theoretical analyses [6, 7], consideration of social network structure and demographics [8], and attempts to elucidate prototypical models combining both mechanisms [9]. A dedicated monograph on behavioural dynamics was published in 2013 [10], and a further review paper in 2015 [11]. Spreading of behaviour continues to inspire mathematical developments, for example development of differential equation systems with cubic and higher-order interaction terms [12], use of multiplex network theory [13], and investigation of paradoxical behaviour [14]. The COVID-19 pandemic, seeing significant impacts from behaviour, also saw attempts to advance this field [15].
Despite this activity, important challenges remain in developing a generalised model of epidemiological and behavioural dynamics, including balancing model complexity, interpretability and capability to validate them with empirical data [16]. And at the same time, few models have attempted to capture specific behaviours, preferring to adopt an intervention-agnostic approach.
1.2 Outline of this study
In this paper, we define and analyse an extension to the Behaviour and Disease (BaD) model of [17, 18] to the behavioural dynamics specific to (symptomatic) testing and isolation. This is because testing and isolation are known to have both a strong behavioural component and a significant impact on disease transmission [19, 20]. In addition, in terms of data availability it is also possible to monitor both test results and compliance with testing policy over time using community surveys [21].
We structure the remainder of the paper as follows. We initially present our model that incorporates dynamic testing behaviour and infectious disease transmission processes (Section 2). We then conduct theoretical analyses of the model (Section 3) to give insight on the early dynamics and endemicity of the system; specifically, we derive its basic reproduction number $\mathcal{R}_{0}$ (Section 3.1), compute its steady states (Section 3.2) and analyse stability criteria (Section 3.3). Note that derivation of quantities such as final size in the absence of waning immunity are not believed to be easily available for this class of models using existing methods (Röst, personal communication relating to the model published in [22]). Accordingly, to investigate how temporal and cumulative outbreak metrics depend on model parameter values we describe the results of numerical simulations (Section 4). We conclude by discussing the findings and their implications (Section 5).
2 Behaviour and Disease (BaD) model for symptomatic testing
In this section we introduce our mathematical model, which extends the BaD model of [17] to incorporate dynamics that model symptomatic testing. The key concept behind this model is the distinction between individuals who do not intend to seek testing when symptomatic, referred to as non-behavers (labelled $N$) and those who seek testing when symptomatic, referred to as behavers (labelled $B$). We emphasise, however, that while we use the nomenclature of behavers and non-behavers for consistency with the other literature, we are considering $B$ individuals as those who hold an intention to test if symptomatic, and so may fail to do this if, for example, there is a lack of testing capacity. This means we parameterise the model in terms of the effectiveness of testing as an intervention, including the real-world factors that can obstruct access to testing, rather than its efficacy under ideal conditions that is primarily determined by test sensitivity.
Our pathogen dynamical framework is built upon the standard SEIRS compartmental framework: susceptible individuals ($S$) become exposed ($E$), progress to an infectious state ($I$, $A$ or $T$ as explained below) and then recover ($R$), with a non-zero rate of waning immunity meaning that recovered individuals can become susceptible again. To account for symptomatic testing behaviour, we further split the infectious class into asymptomatics (or pauci-symptomatics) labelled $A$, and symptomatics who are labelled $I$ if they do not seek or do not receive a positive test, and $T$ if they seek and receive a positive test. The transitions between epidemiological states follow the standard SEIRS framework, while the changes in testing behaviour are governed by the rates of behaviour uptake ($\omega$) and abandonment ($\alpha$). Figure 1 shows the structure of this model.
Let $B(t)$ denote the proportion of the population at time $t$ that would test if showing symptoms, and $N(t)$ denote the proportion that would not. We will henceforth mostly suppress explicit dependence on $t$ for notational compactness. For each state class $X\in\{S,E,I,A,R\}$, $X_{N}$ and $X_{B}$ denote the proportions of non-behavers and behavers, respectively. Additionally, let the proportion of symptomatic individuals who seek testing and test positive be $T$.
We adopt the same functional forms for the rates of behaviour uptake and behaviour loss as the original BaD model [17]. The rate of behaviour uptake is governed by
$\omega(B,T)=\phi_{1}(B)+\phi_{2}(T)+\omega_{3}=\omega_{1}B+\omega_{2}T+\omega_{3}+\mathcal{O}(B,T)^{2}$ (1)
where $\phi_{1}$ represents the influence of social contagion on testing behaviour, $\phi_{2}$ accounts for behaviour uptake driven by the perception of illness threat, and $\omega_{3}$ captures spontaneous adoption of the behaviour. The rate of behaviour loss is
$\alpha(N)=\psi(N)+\alpha_{2}=\alpha_{1}N+\alpha_{2}+\mathcal{O}(N)^{2}$ (2)
where $\psi$ represents the influence of social contagion on not seeking testing and $\alpha_{2}$ captures spontaneous abandonment of the behaviour. Note that while, for a complex contagion model, the quadratic and higher-order terms on the right-hand sides of Equations (1) and (2) might be large and even provide the dominant contribution to $\omega$ and $\alpha$, we will only include linear terms in our analysis when an explicit form is required.
The BaD model for symptomatic testing is described by the following differential equations with the parameters detailed in Table 1.
$\frac{\mbox{d}S_{N}}{\mbox{d}t}$ $=-\lambda S_{N}+\nu R_{N}-\omega(B,T)S_{N}+\alpha(N)S_{B}$ (3)
$\frac{\mbox{d}E_{N}}{\mbox{d}t}$ $=\lambda S_{N}-\sigma E_{N}-\omega(B,T)E_{N}+\alpha(N)E_{B}$ (4)
$\frac{\mbox{d}A_{N}}{\mbox{d}t}$ $=p_{A}\sigma E_{N}-\gamma A_{N}-\omega(B,T)A_{N}+\alpha(N)A_{B}$ (5)
$\frac{\mbox{d}I_{N}}{\mbox{d}t}$ $=\left(1-p_{A}\right)\sigma E_{N}-\gamma I_{N}-\omega(B,T)I_{N}+\alpha(N)I_{B}$ (6)
$\frac{\mbox{d}R_{N}}{\mbox{d}t}$ $=\gamma(A_{N}+I_{N})-\nu R_{N}-\omega(B,T)R_{N}+\alpha(N)R_{B}$ (7)
$\frac{\mbox{d}S_{B}}{\mbox{d}t}$ $=-q_{B}\lambda S_{B}+\nu R_{B}+\omega(B,T)S_{N}-\alpha(N)S_{B}$ (8)
$\frac{\mbox{d}E_{B}}{\mbox{d}t}$ $=q_{B}\lambda S_{B}-\sigma E_{B}+\omega(B,T)E_{N}-\alpha(N)E_{B}$ (9)
$\frac{\mbox{d}A_{B}}{\mbox{d}t}$ $=p_{A}\sigma E_{B}-\gamma A_{B}+\omega(B,T)A_{N}-\alpha(N)A_{B}$ (10)
$\frac{\mbox{d}I_{B}}{\mbox{d}t}$ $=\left(1-p_{A}\right)\left(1-p_{T}\right)\sigma E_{B}-\gamma I_{B}+\omega(B,T)I_{N}-\alpha(N)I_{B}$ (11)
$\frac{\mbox{d}T}{\mbox{d}t}$ $=\left(1-p_{A}\right)p_{T}\sigma E_{B}-\gamma T$ (12)
$\frac{\mbox{d}R_{B}}{\mbox{d}t}$ $=\gamma(A_{B}+I_{B}+T)-\nu R_{B}+\omega(B,T)R_{N}-\alpha(N)R_{B}.$ (13)
The force of infection on non-behavers is given by
$\lambda=\beta\left(I_{N}+I_{B}+q_{A}\left(A_{N}+A_{B}\right)+q_{T}T\right)$
whereas for behavers it is $q_{B}\lambda$.
Figure 1: A compartmental diagram illustrating the BaD model for symptomatic testing. We distinguish between individuals who do not seek testing when symptomatic, referred to as non-behavers (labelled $N$) and those who seek testing when symptomatic, referred to as behavers (labelled $B$). Changes in testing behaviour are governed by the rates of behaviour uptake ($\omega$) and abandonment ($\alpha$). The epidemiological states are susceptible ($S$), exposed ($E$), asymptomatic infectious (or pauci-symptomatics) labelled $A$, symptomatic infectious who are labelled $I$ if they do not seek or do not receive a positive test and $T$ if they seek and receive a positive test, and those who are recovered ($R$). The transitions between epidemiological states follow a standard SEIRS framework, with waning immunity where recovered individuals can become susceptible again. We define the model parameters in Table 1.
Table 1: Behavioural and epidemiological parameters for the model. The parameters and functions defined by Greek letters are in units of $\text{time}^{-1}$, which have been scaled to such that the infectious period is one. All other parameters are dimensionless. Where possible, the last column contains references for the parameter choices. We chose all other parameters to take a priori reasonable values with a non-trivial impact on the dynamics.
Parameter Description Value Ref.
$\lambda(t)$ the force of infection
      $\beta$ the transmission rate $4.151$ [23]
      $q_{A}$ the reduction in infectiousness from being asymptomatic $0.58$ [24]
      $q_{T}$ the reduction in infectiousness from isolation $0.25$
      $q_{B}$ the reduction in susceptibility from the testing behaviour $0.5$ [17]
$\sigma$ the transition rate from $E$ to $I$ (average latency period is $1/\sigma$) $2.5$ [25]
$\gamma$ the recovery rate (average infectious period is $1/\gamma$) $1.0$
$\nu$ the rate of waning immunity (average immune period is $1/\nu$) $0.05$ [26]
$p_{A}$ probability of not developing symptoms $0.18$ [24]
$p_{T}$ test effectiveness for symptomatic individuals $0.9$
$\omega(t)$ the “force of infection" for testing behaviour uptake
      $\omega_{1}$ the social transmission rate $0.25$
      $\omega_{2}$ rate of response to perceived illness threat $7.0$
      $\omega_{3}$ spontaneous uptake rate $0.0125$
$\alpha(t)$ the “force of infection" for testing behaviour abandonment
      $\alpha_{1}$ the social abandonment rate $0.2$
      $\alpha_{2}$ spontaneous abandonment rate $0.1$
For non-behavers, we can express the total proportion as $N=S_{N}+E_{N}+A_{N}+I_{N}+R_{N}.$ Similarly, for behavers, the total proportion is $B=S_{B}+E_{B}+A_{B}+I_{B}+T+R_{B}.$ In the absence of infection ($S_{N}=N$ and $S_{B}=B$) and ignoring higher order terms, the dynamics become
$\frac{\mbox{d}N}{\mbox{d}t}$ $=-\omega(B,0)N+\alpha(N)B$
$=-\left(\omega_{1}B+\omega_{3}\right)N+\left(\alpha_{1}N+\alpha_{2}\right)B$
with $B=1-N$. Linearising about small initial behaviour ($B(0)\ll 1$) we have $N\approx 1$ and, if $\omega_{3}=0$, behaviour spreads through social pressure if $\omega_{1}>\alpha_{1}+\alpha_{2}$. Following [17], we define the social reproduction number to be
$\mathcal{R}_{0}^{B}=\frac{\omega_{1}}{\alpha_{1}+\alpha_{2}}\,.$
The steady state solves
$F(N)=\left(\alpha_{1}-\omega_{1}\right)N^{2}-\left(\alpha_{1}-\alpha_{2}-\omega_{1}-\omega_{3}\right)N-\alpha_{2}=0$ (14)
As $F(0)=-\alpha_{2}<0$ and $F(1)=\omega_{3}>0$, there exists a unique solution $N_{0}\in(0,1)$ with $F(N_{0})=0$. Notice that if $\omega_{3}>0$ then $B_{0}>0$ regardless of choice of $\omega_{1}$, that is, if $\omega_{3}>0$ then $(N,B)=(1,0)$ is not a steady state of the system.
3 Theoretical analysis
3.1 The basic reproduction number
3.1.1 In the absence of behaviour
If $\omega_{3}=0$ and $\mathcal{R}_{0}^{B}<1$ then the infection-free steady state has $S_{N}=1$ and $S_{B}=0$. The epidemic takes off if
$\mathcal{R}_{0}^{D}=\frac{\left(p_{A}q_{A}+1-p_{A}\right)\beta}{\gamma}>1\,.$
We call $\mathcal{R}_{0}^{D}$ the behaviour-free reproduction number. Note the threshold $\mathcal{R}_{0}^{D}$ also holds if $q_{B}=q_{T}=1$, that is, there is no effect on the transmission dynamics from cautious behaviour or isolation.
3.1.2 In the presence of behaviour
We now consider the full model where both behaviour and infection are present (Equations (3)-(13)). Here we calculate the next-generation matrix. The model has two states-at-infection, $E_{N}$ and $E_{B}$. Let $t_{N}(X)$ denote the expected time an individual spends in compartment $X$ following state-at-infection $E_{N}$. Similarly, let $t_{B}(X)$ represent the corresponding time spent in compartment $X$ following state-at-infection $E_{B}$.
Defining $\alpha_{0}=\alpha(N_{0})$ and $\omega_{0}=\omega(B_{0},0)$, we derive the following expressions for the expected time spent in each compartment. For individuals with state-at-infection $E_{N}$, we have
$t_{N}(A_{N})$ $=\frac{p_{A}\left[\alpha_{0}\left(\sigma+\alpha_{0}+\omega_{0}\right)+\gamma\left(\sigma+\alpha_{0}\right)\right]}{\gamma\left(\sigma+\alpha_{0}+\omega_{0}\right)\left(\gamma+\alpha_{0}+\omega_{0}\right)}$ (15)
$t_{N}(I_{N})$ $=\frac{\left(1-p_{A}\right)\left[\alpha_{0}\left(\sigma+\alpha_{0}+\omega_{0}\right)+\gamma\left(\sigma+\alpha_{0}\right)-p_{T}\alpha_{0}\omega_{0}\right]}{\gamma\left(\sigma+\alpha_{0}+\omega_{0}\right)\left(\gamma+\alpha_{0}+\omega_{0}\right)}$
and
$t_{N}(A_{B})$ $=\frac{p_{A}\omega_{0}\left(\gamma+\sigma+\alpha_{0}+\omega_{0}\right)}{\gamma\left(\sigma+\alpha_{0}+\omega_{0}\right)\left(\gamma+\alpha_{0}+\omega_{0}\right)}$ (16)
$t_{N}(I_{B})$ $=\frac{\left(1-p_{A}\right)\omega_{0}\left[\sigma+\alpha_{0}+\left(1-p_{T}\right)\left(\gamma+\omega_{0}\right)\right]}{\gamma\left(\sigma+\alpha_{0}+\omega_{0}\right)\left(\gamma+\alpha_{0}+\omega_{0}\right)}$
$t_{N}(T)$ $=\frac{\left(1-p_{A}\right)p_{T}\omega_{0}}{\gamma\left(\sigma+\alpha_{0}+\omega_{0}\right)}\,.$
Similarly, for individuals with state-at-infection $E_{B}$, we have
$t_{B}(A_{N})$ $=\frac{p_{A}\alpha_{0}\left(\gamma+\sigma+\alpha_{0}+\omega_{0}\right)}{\gamma\left(\sigma+\alpha_{0}+\omega_{0}\right)\left(\gamma+\alpha_{0}+\omega_{0}\right)}$ (17)
$t_{B}(I_{N})$ $=\frac{\left(1-p_{A}\right)\alpha_{0}\left[\gamma+\alpha_{0}+\left(1-p_{T}\right)\left(\sigma+\omega_{0}\right)\right]}{\gamma\left(\sigma+\alpha_{0}+\omega_{0}\right)\left(\gamma+\alpha_{0}+\omega_{0}\right)}$
and
$t_{B}(A_{B})$ $=\frac{p_{A}\left[\omega_{0}\left(\sigma+\alpha_{0}+\omega_{0}\right)+\gamma(\sigma+\omega_{0})\right]}{\gamma\left(\sigma+\alpha_{0}+\omega_{0}\right)\left(\gamma+\alpha_{0}+\omega_{0}\right)}$ (18)
$t_{B}(I_{B})$ $=\frac{\left(1-p_{A}\right)\left[\omega_{0}(\sigma+\alpha_{0}+\omega_{0})+\gamma(\sigma+\omega_{0})-p_{T}(\sigma+\omega_{0})(\gamma+\omega_{0})\right]}{\gamma\left(\sigma+\alpha_{0}+\omega_{0}\right)\left(\gamma+\alpha_{0}+\omega_{0}\right)}$
$t_{B}(T)$ $=\frac{\left(1-p_{A}\right)p_{T}\left(\sigma+\omega_{0}\right)}{\gamma\left(\sigma+\alpha_{0}+\omega_{0}\right)}\,.$
If an individual with state-at-infection $E_{N}$ or $E_{B}$ results in a total (over time) force of infection $\Lambda_{N}$ or $\Lambda_{B}$ respectively on non-behavers $N$, then
$\Lambda_{N}$ $=\beta\left\{t_{N}(I_{N})+t_{N}(I_{B})+q_{A}\left(t_{N}(A_{N})+t_{N}(A_{B})\right)+q_{T}t_{N}(T)\right\}$
and
$\Lambda_{B}$ $=\beta\left\{t_{B}(I_{N})+t_{B}(I_{B})+q_{A}\left(t_{B}(A_{N})+t_{B}(A_{B})\right)+q_{T}t_{B}(T)\right\}.$
The terms $\Lambda_{N}$ and $\Lambda_{B}$ can be further expressed as
$\Lambda_{N}$ $=\frac{\beta}{\gamma}\left\{p_{N}(1-p_{A})+q_{A}p_{A}+(1-p_{N})(1-p_{A})\left[(1-p_{T})+q_{T}p_{T}\right]\right\}$ (19)
$\Lambda_{B}$ $=\frac{\beta}{\gamma}\left\{(1-p_{B})(1-p_{A})+q_{A}p_{A}+p_{B}(1-p_{A})\left[(1-p_{T})+q_{T}p_{T}\right]\right\}$ (20)
where
$p_{N}=\frac{\sigma+\alpha_{0}}{\sigma+\alpha_{0}+\omega_{0}}\qquad\text{and}\qquad p_{B}=\frac{\sigma+\omega_{0}}{\sigma+\alpha_{0}+\omega_{0}}$
are the proportions of time that individual with state-at-infection $E_{N}$ will spend in $N$ before becoming infectious and an individual with state-at-infection $E_{B}$ will spend in $B$ before becoming infectious, respectively.
From Equation (19), we can observe how different factors contribute to the force of infection from individuals with state-at-infection $E_{N}$. The first term $p_{N}(1-p_{A})$ represents the proportion of those individuals who become infectious and symptomatic while in state $N$. The second term $q_{A}p_{A}$ represents the proportion of those who become asymptomatic (note that this term combines contributions from both behaviour states). The expression $(1-p_{N})(1-p_{A})$ in the third and fourth terms describes the proportion of individuals who transition from state $N$ to state $B$ before becoming symptomatic while in $B$; the third term component $(1-p_{T})$ represents those who do not receive a positive test, which could be due to imperfect test sensitivity or to not testing despite having formed the intention to do so, while the fourth term component $q_{T}p_{T}$ represents those who receive a positive test. A similar interpretation applies to Equation (20), describing infections due to those with state-at-infection $E_{B}$.
Finally, the next generation matrix can be written as
$\mathbf{K}=\left(\begin{array}[]{cc}\Lambda_{N}N_{0}&\Lambda_{B}N_{0}\\ q_{B}\Lambda_{N}B_{0}&q_{B}\Lambda_{B}B_{0}\end{array}\right).$
As $\det(\mathbf{K})=0$, the basic reproduction number of the BaD SEIRS model for symptomatic testing is given by
$\mathcal{R}_{0}=\Lambda_{N}N_{0}+q_{B}\Lambda_{B}B_{0}$
where $\Lambda_{N},\Lambda_{B}$ are given by Equation (19) and (20), $N_{0}$ solves Equation (14) and $B_{0}=1-N_{0}.$ Note that $\Lambda_{N}$ and $\Lambda_{B}$ can be rewritten as
$\Lambda_{N}$ $=\mathcal{R}_{0}^{D}-\frac{\beta(1-p_{N})(1-p_{A})(1-q_{T})p_{T}}{\gamma}$
and
$\Lambda_{B}$ $=\mathcal{R}_{0}^{D}-\frac{\beta p_{B}(1-p_{A})(1-q_{T})p_{T}}{\gamma}$
where $\mathcal{R}_{0}^{D}$ is the behaviour-free reproduction number. Writing
$\Delta\mathcal{R}=\frac{\beta(1-p_{A})(1-q_{T})p_{T}}{\gamma}$
shows that
$\mathcal{R}_{0}=\underbrace{N_{0}\mathcal{R}_{0}^{D}}_{\text{(i)}}+\underbrace{q_{B}B_{0}\mathcal{R}_{0}^{D}}_{\text{(ii)}}-\underbrace{(1-p_{N})N_{0}\Delta\mathcal{R}}_{\text{(iii)}}-\underbrace{q_{B}p_{B}B_{0}\Delta\mathcal{R}}_{\text{(iv)}}\,.$ (21)
In the Equation (21): (i) captures the spread of infection from non-isolators (those in $A$ and $I$) to non-behavers (those in $N$); (ii) captures the spread of infection from non-isolators to behavers (those in $B$); (iii) captures the reduction in infections in non-behavers caused by isolation (those in $T$), and; (iv) captures the reduction in infections in behavers caused by isolation. Note that in the limiting case where $B_{0}$ approaches 00, $p_{N}$ approaches $1$ and Equation (21) shows that $\mathcal{R}_{0}$ approaches $\mathcal{R}_{0}^{D}$.
3.1.3 The effective reproduction number
The effective reproduction number can be defined using Equation (21) such that
$\mathcal{R}(t)=S_{N}(t)\mathcal{R}_{0}^{D}+q_{B}S_{B}(t)\mathcal{R}_{0}^{D}-(1-p_{N}(t))S_{N}(t)\Delta\mathcal{R}-q_{B}p_{B}(t)S_{B}(t)\Delta\mathcal{R}$ (22)
where
$p_{N}(t)=\frac{\sigma+\alpha(N(t))}{\sigma+\alpha(N(t))+\omega(B(t),T(t))}$ and $p_{B}(t)=\frac{\sigma+\omega(B(t),T(t))}{\sigma+\alpha(N(t))+\omega(B(t),T(t))}\,.$
It is important to note that $\mathcal{R}(t)$ is initial condition dependant. For example, when the system starts near equilibrium such that $S_{N}(0)\approx N_{0}$ and $S_{B}(0)\approx B_{0}$, we observe that $\mathcal{R}(0)\approx\mathcal{R}_{0}$.
Alternatively, consider where $S_{N}(0)\approx 1$ and $S_{B}(0)$ is negligible. This captures situations where, for example, test availability may be extremely limited in early stage of an epidemic or the importance of testing is not widely known for a new pathogen. For these conditions, we observe that
$\mathcal{R}(0)\approx\mathcal{R}_{0}^{D}-(1-p_{N}(0))\Delta\mathcal{R}$
where
$1-p_{N}(0)\approx\frac{\omega_{3}}{\sigma+\alpha_{1}+\alpha_{2}+\omega_{3}}$
captures the reduction in the behaviour-free reproduction number due to a small number of people spontaneously testing in the early stage of the epidemic.
3.2 The infected steady states
Let $X^{*}$ denote the steady state for compartment $X$. When infection is present, the dynamics of behaviour are governed by
$\frac{\mbox{d}N}{\mbox{d}t}$ $=-\omega(B,T)N+\alpha(N)(B-T)\,.$ (23)
Note that there exists a steady state solution $(N^{*},B^{*})$ with $B^{*}=1-N^{*}$ that depends on the steady state $T^{*}$. That is, $N^{*}$ solves the quadratic
$G(N)=(\omega_{1}-\alpha_{1})N^{2}-(\omega_{2}T^{*}+\omega_{1}+\omega_{3}+\alpha_{2}-\alpha_{1}(1-T^{*}))N+\alpha_{2}(1-T^{*})=0.$
We see that $G(0)=\alpha_{2}(1-T^{*})>0$ and $G(1)=-[(\omega_{2}+\alpha_{1}+\alpha_{2})T^{*}+\omega_{3}]<0$. Hence, given $T^{*}$ there exists a unique $N^{*}\in(0,1)$ with $G(N^{*})=0$.
Now, fix a value for the prevalence of symptomatic cases $O^{*}=I_{N}^{*}+I_{B}^{*}+T^{*}$. Then
$E^{*}$ $=E_{N}^{*}+E_{B}^{*}=\frac{\gamma}{(1-p_{A})\sigma}O^{*}$ (24)
$A^{*}$ $=A_{N}^{*}+A_{B}^{*}=\frac{p_{A}}{(1-p_{A})}O^{*}$ (25)
$R^{*}$ $=R_{N}^{*}+R_{B}^{*}=\frac{\gamma}{\nu(1-p_{A})}O^{*}$ (26)
$S^{*}$ $=S_{N}^{*}+S_{B}^{*}=1-(E^{*}+A^{*}+O^{*}+R^{*})\,.$ (27)
Also, knowing $T^{*}$ fixes $N^{*}$, $B^{*}=1-N^{*}$, and
$\lambda^{*}$ $=\beta\left(I^{*}+q_{A}A^{*}+q_{T}T^{*}\right)$ (28)
$\omega^{*}$ $=\omega_{1}B^{*}+\omega_{2}T^{*}+\omega_{3}$ (29)
$\alpha^{*}$ $=\alpha_{1}N^{*}+\alpha_{2}$ (30)
where $I^{*}=I_{N}^{*}+I_{B}^{*}$. Now, setting Equations (4) and (9) to zero gives
$\lambda^{*}S_{N}^{*}$ $=(\omega^{*}+\sigma)E^{*}-(\alpha^{*}+\omega^{*}+\sigma)E_{B}^{*}$ (31)
$\lambda^{*}S_{B}^{*}$ $=\frac{-\omega^{*}E^{*}+(\alpha^{*}+\omega^{*}+\sigma)E_{B}^{*}}{q_{B}}\,.$ (32)
Adding Equations (31) and (32) gives
$E_{B}^{*}=\frac{q_{B}\lambda^{*}S^{*}+(\omega^{*}-q_{B}(\omega^{*}+\sigma))E^{*}}{(1-q_{B})(\alpha^{*}+\omega^{*}+\sigma)}\,.$ (33)
Setting Equations (10) and (11) to zero respectively gives
$A_{B}^{*}$ $=\frac{p_{A}\sigma E_{B}^{*}+\omega^{*}A^{*}}{\gamma+\alpha^{*}+\omega^{*}}$ (34)
$I_{B}^{*}$ $=\frac{(1-p_{A})(1-p_{T})\sigma E_{B}^{*}+\omega^{*}I^{*}}{\gamma+\alpha^{*}+\omega^{*}}\,.$ (35)
Next, setting Equation (13) to zero gives
$R_{B}^{*}=\frac{\gamma(A_{B}^{*}+I_{B}^{*}+T^{*})+\omega^{*}R^{*}}{\alpha^{*}+\omega^{*}+\nu}\,.$ (36)
Thus, given $O^{*}$ and $T^{*}$, the steady states are given by
$S_{N}^{*}$ $=S^{*}-S_{B}^{*}$ $S_{B}^{*}$ $=B^{*}-(E_{B}^{*}+A_{B}^{*}+I_{B}^{*}+T^{*}+R_{B}^{*})$
[…]
Finally, setting Equations (3) and (12) to zero gives the simultaneous equations
$T^{*}$ $=\frac{(1-p_{A})p_{T}\sigma}{\gamma}E_{B}^{*}$
$(\lambda^{*}+\omega^{*})S_{N}^{*}$ $=\alpha^{*}S_{B}^{*}+\nu R_{N}^{*}$
which fixes $I^{*}$ and $T^{*}$ (and hence $O^{*}$).
3.3 Stability
To analyse the stability of the system, we begin by expressing the model in a more compact form. Recall that the evolution of the non-behaver population ($N$) follows
$\frac{\mbox{d}N}{\mbox{d}t}=-\omega(B,T)N+\alpha(N)\left(B-T\right)$ (37)
where $B=1-N$. The dynamics of the non-behavers are given by
$\frac{\mbox{d}S_{N}}{\mbox{d}t}$ $=-\lambda S_{N}-\omega(B,T)S_{N}+\alpha(N)S_{B}+\nu R_{N}$ (38)
$\frac{\mbox{d}E_{N}}{\mbox{d}t}$ $=\lambda S_{N}-\sigma E_{N}-\omega(B,T)E_{N}+\alpha(N)E_{B}$
$\frac{\mbox{d}A_{N}}{\mbox{d}t}$ $=p_{A}\sigma E_{N}-\gamma A_{N}-\omega(B,T)A_{N}+\alpha(N)A_{B}$
$\frac{\mbox{d}I_{N}}{\mbox{d}t}$ $=\left(1-p_{A}\right)\sigma E_{N}-\gamma I_{N}-\omega(B,T)I_{N}+\alpha(N)I_{B}$
where $R_{N}=N-S_{N}-E_{N}-A_{N}-I_{N}$. Similarly, the equations governing the behavers are
$\frac{\mbox{d}S_{B}}{\mbox{d}t}$ $=-q_{B}\lambda S_{B}+\omega(B,T)S_{N}-\alpha(N)S_{B}+\nu R_{B}$ (39)
$\frac{\mbox{d}E_{B}}{\mbox{d}t}$ $=q_{B}\lambda S_{B}-\sigma E_{B}+\omega(B,T)E_{N}-\alpha(N)E_{B}$
$\frac{\mbox{d}A_{B}}{\mbox{d}t}$ $=p_{A}\sigma E_{B}-\gamma A_{B}+\omega(B,T)A_{N}-\alpha(N)A_{B}$
$\frac{\mbox{d}I_{B}}{\mbox{d}t}$ $=\left(1-p_{A}\right)\left(1-p_{T}\right)\sigma E_{B}-\gamma I_{B}+\omega(B,T)I_{N}-\alpha(N)I_{B}$
$\frac{\mbox{d}T}{\mbox{d}t}$ $=\left(1-p_{A}\right)p_{T}\sigma E_{B}-\gamma T$
where $R_{B}=B-S_{B}-E_{B}-A_{B}-I_{B}-T$.
To obtain a matrix representation, we define the following matrices and vectors for non-behavers:
$\mathbf{M}_{N}=\left(\begin{array}[]{cccc}-\lambda-\nu&-\nu&-\nu&-\nu\\ \lambda&-\sigma&0&0\\ 0&p_{A}\sigma&-\gamma&0\\ 0&\left(1-p_{A}\right)\sigma&0&-\gamma\end{array}\right),\qquad\mathbf{n}=\left(\begin{array}[]{c}S_{N}\\ E_{N}\\ A_{N}\\ I_{N}\end{array}\right)$
and for the behavers:
$\mathbf{M}_{B}=\left(\begin{array}[]{cccc}-q_{B}\lambda-\nu&-\nu&-\nu&-\nu\\ q_{B}\lambda&-\sigma&0&0\\ 0&p_{A}\sigma&-\gamma&0\\ 0&\left(1-p_{A}\right)\left(1-p_{T}\right)\sigma&0&-\gamma\end{array}\right),\qquad\mathbf{b}=\left(\begin{array}[]{c}S_{B}\\ E_{B}\\ A_{B}\\ I_{B}\end{array}\right)$
with
$\lambda=\beta\left[I_{N}+I_{B}+q_{A}\left(A_{N}+A_{B}\right)+q_{T}T\right].$
Then the model can be rewritten in matrix form as
$\frac{\mbox{d}N}{\mbox{d}t}$ $=-\omega(B,T)N+\alpha(N)\left(1-N-T\right)$ (40)
$\frac{\mbox{d}T}{\mbox{d}t}$ $=\left(1-p_{A}\right)p_{T}\sigma E_{B}-\gamma T$
with
$\frac{\mbox{d}\mathbf{n}}{\mbox{d}t}$ $=\mathbf{M}_{N}\mathbf{n}-\omega(B,T)\mathbf{n}+\alpha(N)\mathbf{b}+\nu N\mathbf{u}_{1}$ (41)
$\frac{\mbox{d}\mathbf{b}}{\mbox{d}t}$ $=\mathbf{M}_{B}\mathbf{b}+\omega(B,T)\mathbf{n}-\alpha(N)\mathbf{b}+\nu\left(1-N-T\right)\mathbf{u}_{1}$
where $\mathbf{u}_{1}$ is the first standard basis vector, having a first element equal to $1$ and all other elements equal to 00. The Jacobian matrix for this system is structured as
$\mathbf{J}=\left(\begin{array}[]{c|cccc|cccc}&0&0&0&0&0&0&0&0\\ \mathbf{J}^{0}&&&&&&&&\\ &0&0&0&0&0&\left(1-p_{A}\right)p_{T}\sigma&0&0\\ \hline\cr&&&&&&&&\\ \mathbf{J}^{n0}&&&\mathbf{J}^{nn}&&&\hskip 28.45274pt\mathbf{J}^{nb}&&\\ &&&&&&&&\\ \hline\cr&&&&&&&&\\ \mathbf{J}^{b0}&&&\mathbf{J}^{bn}&&&\hskip 28.45274pt\mathbf{J}^{bb}&&\\ &&&&&&&&\\ \end{array}\right)$
where $\mathbf{J}^{0}$ is a $2\times 2$ matrix given by
$\mathbf{J}^{0}=\left(\begin{array}[]{cc}\dfrac{\partial\omega}{\partial B}N+\dfrac{\partial\alpha}{\partial N}\left(1-N-T\right)-\omega-\alpha&-\dfrac{\partial\omega}{\partial T}N-\alpha\\[8.5359pt] 0&-\gamma\end{array}\right)$
$\mathbf{J}^{n0}$ and $\mathbf{J}^{b0}$ are $4\times 2$ matrices given by
$\mathbf{J}^{n0}=\left(\begin{array}[]{cc}\dfrac{\partial\omega}{\partial B}S_{N}+\dfrac{\partial\alpha}{\partial N}S_{B}+\nu&-q_{T}\beta S_{N}-\dfrac{\partial\omega}{\partial T}S_{N}\\[8.5359pt] \dfrac{\partial\omega}{\partial B}E_{N}+\dfrac{\partial\alpha}{\partial N}E_{B}&q_{T}\beta S_{N}-\dfrac{\partial\omega}{\partial T}E_{N}\\[8.5359pt] \dfrac{\partial\omega}{\partial B}A_{N}+\dfrac{\partial\alpha}{\partial N}A_{B}&-\dfrac{\partial\omega}{\partial T}A_{N}\\[8.5359pt] \dfrac{\partial\omega}{\partial B}I_{N}+\dfrac{\partial\alpha}{\partial N}I_{B}&-\dfrac{\partial\omega}{\partial T}I_{N}\end{array}\right),\quad\mathbf{J}^{b0}=\left(\begin{array}[]{cc}-\dfrac{\partial\omega}{\partial B}S_{N}-\dfrac{\partial\alpha}{\partial N}S_{B}-\nu&-q_{B}q_{T}\beta S_{B}+\dfrac{\partial\omega}{\partial T}S_{N}-\nu\\[8.5359pt] -\dfrac{\partial\omega}{\partial B}E_{N}-\dfrac{\partial\alpha}{\partial N}E_{B}&q_{B}q_{T}\beta S_{B}+\dfrac{\partial\omega}{\partial T}E_{N}\\[8.5359pt] -\dfrac{\partial\omega}{\partial B}A_{N}-\dfrac{\partial\alpha}{\partial N}A_{B}&\dfrac{\partial\omega}{\partial T}A_{N}\\[8.5359pt] -\dfrac{\partial\omega}{\partial B}I_{N}-\dfrac{\partial\alpha}{\partial N}I_{B}&\dfrac{\partial\omega}{\partial T}I_{N}\end{array}\right)$
and the other four sub-matrices $\mathbf{J}^{nn}$, $\mathbf{J}^{nb}$, $\mathbf{J}^{bn}$ and $\mathbf{J}^{bb}$ have dimension $4\times 4$, given by
$\mathbf{J}^{nn}=\left(\begin{array}[]{cccc}-\lambda-\omega-\nu&-\nu&-q_{A}\beta S_{N}-\nu&-\beta S_{N}-\nu\\ \lambda&-\sigma-\omega&q_{A}\beta S_{N}&\beta S_{N}\\ 0&p_{A}\sigma&-\gamma-\omega&0\\ 0&\left(1-p_{A}\right)\sigma&0&-\gamma-\omega\end{array}\right)$
$\mathbf{J}^{nb}=\left(\begin{array}[]{cccc}\alpha&0&-q_{A}\beta S_{N}&-\beta S_{N}\\ 0&\alpha&q_{A}\beta S_{N}&\beta S_{N}\\ 0&0&\alpha&0\\ 0&0&0&\alpha\end{array}\right)$
$\mathbf{J}^{bn}=\left(\begin{array}[]{cccc}\omega&0&-q_{A}q_{B}\beta S_{B}&-q_{B}\beta S_{B}\\ 0&\omega&q_{A}q_{B}\beta S_{B}&q_{B}\beta S_{B}\\ 0&0&\omega&0\\ 0&0&0&\omega\end{array}\right)$
$\mathbf{J}^{bb}=\left(\begin{array}[]{cccc}-q_{B}\lambda-\alpha-\nu&-\nu&-q_{A}q_{B}\beta S_{B}-\nu&-q_{B}\beta S_{B}-\nu\\ q_{B}\lambda&-\sigma-\alpha&q_{A}q_{B}\beta S_{B}&q_{B}\beta S_{B}\\ 0&p_{A}\sigma&-\gamma-\alpha&0\\ 0&\left(1-p_{A}\right)\left(1-p_{T}\right)\sigma&0&-\gamma-\alpha\end{array}\right)\,.$
We now assess the stability of the infection-free steady state, where $S_{N}=N_{0}$ and $S_{B}=B_{0}$. The sub-matrices of the Jacobian matrix $J$, evaluated at this equilibrium, are given by
$\mathbf{J}^{0}=\left(\begin{array}[]{cc}\omega_{1}N_{0}+\alpha_{1}B_{0}-\omega(B_{0},0)-\alpha(N_{0})&-\omega_{2}N_{0}-\alpha(N_{0})\\ 0&-\gamma\end{array}\right)$
$\mathbf{J}^{n0}=\left(\begin{array}[]{cc}\omega_{1}N_{0}+\alpha_{1}B_{0}+\nu&-q_{T}\beta N_{0}-\omega_{2}N_{0}\\ 0&q_{T}\beta N_{0}\\ 0&0\\ 0&0\end{array}\right),\quad\mathbf{J}^{b0}=\left(\begin{array}[]{cc}-\omega_{1}N_{0}-\alpha_{1}B_{0}-\nu&-q_{B}q_{T}\beta B_{0}+\omega_{2}N_{0}-\nu\\ 0&q_{B}q_{T}\beta B_{0}\\ 0&0\\ 0&0\end{array}\right)$
$\mathbf{J}^{nn}=\left(\begin{array}[]{cccc}-\omega(B_{0},0)-\nu&-\nu&-q_{A}\beta N_{0}-\nu&-\beta N_{0}-\nu\\ 0&-\sigma-\omega(B_{0},0)&q_{A}\beta N_{0}&\beta N_{0}\\ 0&p_{A}\sigma&-\gamma-\omega(B_{0},0)&0\\ 0&\left(1-p_{A}\right)\sigma&0&-\gamma-\omega(B_{0},0)\end{array}\right)$
$\mathbf{J}^{nb}=\left(\begin{array}[]{cccc}\alpha(N_{0})&0&-q_{A}\beta N_{0}&-\beta N_{0}\\ 0&\alpha(N_{0})&q_{A}\beta N_{0}&\beta N_{0}\\ 0&0&\alpha(N_{0})&0\\ 0&0&0&\alpha(N_{0})\end{array}\right)$
$\mathbf{J}^{bn}=\left(\begin{array}[]{cccc}\omega(B_{0},0)&0&-q_{A}q_{B}\beta B_{0}&-q_{B}\beta B_{0}\\ 0&\omega(B_{0},0)&q_{A}q_{B}\beta B_{0}&q_{B}\beta B_{0}\\ 0&0&\omega(B_{0},0)&0\\ 0&0&0&\omega(B_{0},0)\end{array}\right)$
and
$\mathbf{J}^{bb}=\left(\begin{array}[]{cccc}-\alpha(N_{0})-\nu&-\nu&-q_{A}q_{B}\beta B_{0}-\nu&-q_{B}\beta B_{0}-\nu\\ 0&-\sigma-\alpha(N_{0})&q_{A}q_{B}\beta B_{0}&q_{B}\beta B_{0}\\ 0&p_{A}\sigma&-\gamma-\alpha(N_{0})&0\\ 0&\left(1-p_{A}\right)\left(1-p_{T}\right)\sigma&0&-\gamma-\alpha(N_{0})\end{array}\right).$
where $B_{0}=1-N_{0}$, $\omega(B_{0},0)=\omega_{1}B_{0}+\omega_{3}$ and $\alpha(N_{0})=\alpha_{1}N_{0}+\alpha_{2}$. The infection-free steady state where $S_{N}=N_{0}$ and $S_{B}=B_{0}$ follows standard results for disease dynamics without a behavioural component and so is stable when $\mathcal{R}_{0}<1$.
4 Numerical investigations
Having obtained analytical insights on the early outbreak dynamics and stability criterion of the system, we next apply numerical methods to gather epidemiological and behavioural insights in epidemic and endemic infection regimes.
For the epidemic regime ($\nu=0$), we explore numerically the impact of key model parameters on the dynamics without waning immunity (SEIR-like infection structure), through exploration of the final size, peaks in testing behaviour, and phase diagrams (Section 4.1). For the endemic regime ($\nu>0$), we use a model with waning immunity (SEIRS-like infection structure) to explore the impacts on the steady states and their stability (Section 4.2).
Given our focus on the uptake of willingness to test and isolate, the key model parameters we vary are the testing effectiveness ($p_{T}$), isolation effectiveness ($1-q_{T}$), initial proportion of the population willing to test and isolate ($B(0)$) by varying spontaneous uptake ($\omega_{3}$), the behaviour-free reproduction number ($\mathcal{R}_{0}^{D}$) by varying the transmission rate ($\beta$), and the social reproduction number ($\mathcal{R}_{0}^{B}$) by varying the social transmission rate ($\omega_{1}$). We further explore the impacts of changing each aspect of the uptake rate on the transient dynamics through phase diagrams without waning immunity.
We list the baseline parameters for our numerical investigation in Table 1. We chose the transmission rate, asymptomatic rate, latent period, and immune period to represent a COVID-like illness. Specifically, we set $\beta$ such that the basic reproduction number is $\mathcal{R}_{0}=3.28$[23], we fixed $p_{A}$ and $q_{A}$ at $0.18$ and $0.58$ respectively [24], and fixed $\sigma$ and $\nu$ at $2.5$ [25] and $0.05$ [26] respectively to correspond to a scaled infectious period where $\gamma=1$ representing an average infection period of $10$ days [25]. We selected the behavioural parameters making up $\alpha$ and $\omega$ to replicate waves in testing behaviour observed over multiple SARS-CoV-2 variants in real data [21]. We fixed the reduction in susceptibility due to risk aversion at $q_{B}=0.5$, similarly to previous behaviour and disease models [17]. Lastly, we fixed the parameters $p_{T}$ and $q_{T}$ at $0.9$ and $0.25$ respectively at baseline, and varied these throughout the numerical explorations.
We ran all numerical simulations with initial conditions of $S_{B}(0)=B_{0}$, $I_{N}(0)=10^{-6}$, $S_{N}(0)=1-S_{B}(0)-I_{N}(0)$ and all other compartments empty, where $B_{0}$ is the disease-free equilibrium. The exception to this was when exploring the transitory dynamics without waning immunity, when we instead set $S_{B}(0)=10^{-6}$. This choice was made to ensure the effective reproduction number at time zero was approximately comparable between simulations when investigating the phase diagrams.
We developed and ran code using Python 3.10.12. The code associated with this study is available at https://github.com/Matthew-Ryan1995/BaD_testing_and_isolation.
4.1 Epidemic regime with no waning immunity
Here we report the impact of key model parameters on the dynamics without waning immunity (i.e. a single-epidemic regime, with $\nu=0$), through exploration of the symptomatic final size, peaks in testing behaviour and the effect of behavioural uptake rates on the epidemiological-behavioural dynamics.
4.1.1 Symptomatic final size
The testing parameters (effectiveness of testing, $p_{T}$, and effectiveness of isolation, $1-q_{T}$) exhibit distinct influences over the observed ($T$) and true symptomatic ($O=I+T$) final sizes of the epidemic (Figure 2, rows one and two). Increasing the effectiveness of the test increases the proportion of the observed final size and interacts non-linearly with the innate infectiousness of the pathogen, $\mathcal{R}_{0}^{D}$. Specifically, even with a completely effective test ($p_{T}=1$), if $\mathcal{R}_{0}^{D}$ is small enough we will observe a negligible final size, whereas the true impact of the disease can be significant (Figure 2, (a), (d), (g)). Unsurprisingly, as the effectiveness of the test increases (i.e. higher values of $p_{T}$) and we observe a larger final size, the unobserved final size (Figure 2 (g)) decreases. The effect of isolation in reducing infection ($q_{T}$: column two) has an opposite impact on the final size. Although the effect is not strong for these parameter values, increasing the effect of isolation (by reducing $q_{T}$) can reduce the observed final size (Figure 2 (b)) and the true symptomatic final size (Figure 2 (e)). This relationship is driven by less contact with the susceptible community as isolation improves.
Initial behaviour in the population ($B(0)$; column three of Figure 2) has a strong and non-linear effect on the symptomatic final size of the epidemic, both observed and undetected. Increasing the initial proportion of the population willing to test and isolate at the start of the epidemic can reduce the final size to a negligible value by driving the reproduction number $\mathcal{R}_{0}$ down. For example, an initial proportion of the population willing to test and isolate exceeding $60\%$ can suppress an outbreak when the behaviour-free reproduction number $\mathcal{R}_{0}^{D}$ is up to $2.1$.
These results suggest that having a population primed to be willing to test and isolate may have a substantial influence in controlling a new epidemic. For example, if a new pathogen has been observed in a different population, efforts to increase test availability and testing willingness in the population may provide a strong preparedness plan. A strong influence of the feedback between perception of illness threat and final size is also observed (Figure 2, (c), (f)). Specifically, for a more infectious disease ($\mathcal{R}_{0}^{D}>4$), increasing the initial populace willing to test and isolate increases the amount of infection seen throughout the epidemic as $T$ increases. However, the actual symptomatic final size ($O$) is reduced due to the behavioural feedback mechanism.
Infection peaks show qualitatively similar patterns (Supplementary Figure S2).
Figure 2: Impact of test effectiveness (${p_{T}}$), isolation effectiveness ($1-{q_{T}}$), initial condition of behaviour (${B(0)}$) and the behaviour-free reproduction number (${\mathcal{R}_{0}^{D}}$) on observed symptomatic (${T}$), undetected symptomatic (${I}$), and total symptomatic (${O=I+T}$) final sizes. The colour bars represent the final size as a percentage of the population and are consistent across each row. Note, we vary the behaviour-free reproduction number instead of $\mathcal{R}_{0}$ due to the dependence of $\mathcal{R}_{0}$ on the vertical axes. The dashed grey lines show $\mathcal{R}_{0}=1$ and the black cross indicates the baseline parameter values from Table 1.
4.1.2 Peaks in testing behaviour
The peak of behaviour over the epidemic describes the maximum proportion of the population at a single time point willing to test and isolate if showing symptoms (Figure 3). Given the feedback mechanism between perception of illness threat and testing behaviour, unsurprisingly increasing the infectiousness of the pathogen of interest ($\mathcal{R}_{0}^{D}$) increases the peak behaviour observed over the epidemic. Changing the initial population willing to test and isolate ($B(0)$) and the social reproduction number ($\mathcal{R}_{0}^{B}$) have similar qualitative effects on peak behaviour (Figure 3, (a), (b)). When the basic reproduction number $\mathcal{R}_{0}<1$ (grey dashed line), we observe the peak of behaviour is kept constant at the disease-free equilibrium $B_{0}$. When $\mathcal{R}_{0}>1$, we observe a ‘waterfall’ like contour where peak behaviour rapidly increases with the behaviour-free reproduction number $\mathcal{R}_{0}^{D}$. This behaviour is driven by the perception of illness threat $\omega_{2}$ and is more pronounced when altering $\omega_{3}$ through $B(0)$ than when altering $\omega_{1}$ through $\mathcal{R}_{0}^{B}$. A notable and interesting relationship occurs with the effectiveness of the test $p_{T}$ (Figure 3 (c)), which shows a non-linear relationship between pathogen infectiousness and increased behaviour. Since the feedback mechanism in the model depends on the observed number of cases $T$, we only observe an increase in the peak behaviour from this mechanism when the testing effectiveness is sufficiently large. Finally, when altering the testing parameters $p_{T}$ and $q_{T}$, observe the visual correlation between peak behaviour (Figure 3, (c), (d)) and detected cases (Figure 2, (a), (b)). This correlation underpins the relationship between behaviour and testing through the perception of illness threat feedback mechanism.
Figure 3: The impact of key model parameters on the peak population willing to test (${B}$) for different behaviour-free reproduction numbers (${\mathcal{R}_{0}^{D}}$). (a) The initial proportion willing to test and isolate ($B(0)$) varies through changing spontaneous uptake ($\omega_{3}$); (b) test effectiveness ($p_{T}$) varies; (c) isolation effectiveness ($1-q_{T}$) varies, and; (d) the social reproduction number ($R_{0}^{B}$) varies through changing the social influence of behaviour ($\omega_{1}$). Note we vary the behaviour-free reproduction number instead of $\mathcal{R}_{0}$ because of the dependence of $\mathcal{R}_{0}$ on the vertical axes. The dashed grey line shows $\mathcal{R}_{0}=1$ and the black cross shows the baseline parameter values from Table 1.
4.1.3 Effect of behavioural uptake rates on the dynamics
The transitory dynamics of the BaD testing and isolation model vary based on the behavioural uptake parameters (Figure 4). Across all three uptake parameters, the observed epidemic (Figure 4, top row) is substantially smaller than the true epidemic in the population (Figure 4, middle row). Two points of particular note stand out across all three behavioural constructs. First, increasing uptake of behaviour makes the epidemic look more severe through an increased reporting of cases ($T$) but the true effect of the epidemic is reduced ($O+A$). This is particularly evident when increasing the perception of illness threat (Figure 4 (b)), which is driven by the feedback mechanism linking $T$ and $B$. Here, we observe a “reporting epidemic” that results in a larger behavioural response from the population, reducing the overall impact of the epidemic. Second, removing any individual construct is sufficient to make sure the epidemic is hardly observed (Figure 4, dashed lines). This suggests an important interplay between the three behavioural constructs in ensuring that individuals in the population test and isolate in response to a new infection. However, each uptake mechanism has a different effect on the observed and true epidemic.
Social uptake of behaviour ($\omega_{1}$) can control the epidemic (Figure 4, (a), (d)). Increasing social pressure has the effect that we observe a larger epidemic, but the peak of the true epidemic is reduced due to increased public awareness and isolation. When social pressure is increased sufficiently, we obtain a greatly reduced overall epidemic and observe almost every case. This effect is driven by increasing $\mathcal{R}_{0}^{B}>1$ so that the behaviour “takes off” socially, leading to a large and quick uptake of testing behaviour. This social spread of behaviour also ensures that the disease-free equilibrium ($B_{0}$) is such that the infection cannot spread easily in the population (Figure 4 (g)).
Spontaneous uptake of behaviour ($\omega_{3}$) has a similar effect as social pressure (Figure 4, (c), (f)). Increasing spontaneous uptake in the model reduces the true epidemic size while increasing the peak observed number of cases. Removal of spontaneous uptake ensures that the epidemic is not observed in this model. Due to the baseline parameter values (Table 1), without a small influence of spontaneous behavioural uptake there is not enough awareness of testing in the population for testing to “take off”, leading to a severe lack of testing and reporting of the epidemic. In this setting, we would observe a silent epidemic where many individuals become unwell but do not seek testing and isolation. Increasing spontaneous uptake sufficiently has the same effect in controlling the epidemic as social pressure. Like social pressure, this control of the epidemic is driven by the speed of behaviour uptake and the increase in $B(0)$ (Figure 4 (i)).
In contrast, the perception of illness threat ($\omega_{2}$) cannot control the epidemic (Figure 4, (b), (e)). When perception of illness threat is increased, we observe a ramp up of reported cases followed by a sharp decline and conclusion of the epidemic. This is driven by the feedback mechanism in the model: the more cases reported to the population through the $T$ compartment, the sharper uptake of testing behaviour. This feedback mechanism is the reason that perception of illness threat cannot control the epidemic. Since $\omega_{2}$ does not influence the basic reproduction number or the disease free equilibrium (Figure 4 (h)), for the feedback mechanism to take effect we need to observe cases of infection in the population.
The effect of behavioural parameters on the transient dynamics are similar in an endemic regime, albeit with trajectories converging to steady states (Figure S3).
Figure 4: Phase diagrams for different behavioural parameters fixing ${\mathcal{R}_{0}^{D}=3.28}$ in the epidemic regime. Column one corresponds to changes in social influence ($\omega_{1}$), column two corresponds to changes in perception of illness threat ($\omega_{2}$), and column three corresponds to spontaneous uptake ($\omega_{3}$), respectively. The rows show the susceptible versus observed epidemic ($S$ vs. $T$), the susceptible versus true epidemic ($S$ vs. $O+A$), and the behaviour versus observed epidemic ($B$ vs. $T$) phase planes. The dashed line represents near removal of the behaviour construct, the dotted lines represent a reduction of the baseline value, the solid line represents the baseline values, and the dash-dot lines represent an increase in the baseline values. The solid lines all represent the baseline values from Table 1, ensuring the solid lines are comparable across each column. Each simulation was run with the initial conditions $S_{B}(0)=10^{-6},I_{N}(0)=10^{-6},S_{N}=1-S_{B}-I_{N}$ and all other compartments empty: this ensured the early-stage dynamics of each model were approximately comparable.
4.2 Endemic regime with waning immunity
Here we detail our findings for an endemic regime with $\nu>0$ as detailed in Table 1. We report on the impact of testing and isolation effects on endemic steady states, the impact of behaviour and disease contagiousness on endemic states, and the stability of steady states.
4.2.1 Testing and isolation effects on endemic steady states
Testing effectiveness ($p_{T}$) and isolation effectiveness ($1-q_{T}$) can have a non-linear effect on the endemic steady states (Figure 5). However, similarly to the effects on the final size in the epidemic regime (Figure 2), for the steady state infection prevalence in the endemic regime improved testing effectiveness $p_{T}$ also has the more substantial effect. We observe the largest reduction in endemic infection prevalence (Figure 5, (a), (b)) for both high test effectiveness (large $p_{T}$) and effective isolation (low $q_{T}$; high $1-q_{T}$); this combination is also linked to the largest reduction in $\mathcal{R}_{0}$ (Figure 5 (f)). For lower levels of testing effectiveness, isolation has a negligible effect on the observed and unobserved symptomatic infection rates and the proportion of the population willing to test and isolate if symptomatic (Figure 5, (c), (d), (e)). When testing effectiveness is high, we note a reduction in observed and unobserved symptomatic infection prevalences for improved isolation. In these settings, the improved isolation is contributing to an overall reduction in infection, so observed cases are similarly reduced. Due to the reduced number of cases observed for high values of $p_{T}$ and low values of $q_{T}$ (high $1-q_{T}$), we observe a corresponding reduction in the proportion of the population willing to test and isolate for the infection in the steady state of the system (Figure 5 (e)).
Figure 5: The relationship between isolation & test effectiveness and endemic states. In each panel we display outputs for different combinations of isolation effectiveness ($1-q_{T}$, y-axis) and test effectiveness ($p_{T}$, x-axis). In all simulations we used a fixed behaviour-free reproduction number of $\mathcal{R}_{0}^{D}=3.28$. (a) Total endemic infection state in the population ($O+A$); (b) Symptomatic endemic infection state in the population ($O$); (c) Observed symptomatic endemic infection state in the population ($T$); (d) Undetected symptomatic infection endemic state in the population ($I$); (e) Proportion of the population willing to test and isolate if symptomatic in the steady state of the population ($B$); (f) Basic reproduction number ($\mathcal{R}_{0}$). For (a)-(e), colour bars show the percentage of the population in each state. For (f), the colour bar shows the change in reproduction number.
4.2.2 Behaviour and disease contagiousness on endemic states
Both the innate infectiousness of the pathogen ($\mathcal{R}_{0}^{D}$) and the social contagiousness of the testing behaviour ($\mathcal{R}_{0}^{B}$) have significant impacts on reported and unreported case numbers in the endemic steady state (Figure 6). The boundary for elimination of the infection by reducing $\mathcal{R}_{0}<1$ is non-linear in the behaviour and diseases contagiousness measures (Figure 6 (f)). This non-linearity suggests that for a more infectious disease (larger $\mathcal{R}_{0}^{D}$) we require a substantial increase in the social influence of testing behaviour (larger $\mathcal{R}_{0}^{B}$) to eliminate the infection in the steady state. This corresponds to a larger increase in the percentage of the population willing to test and isolate if symptomatic (Figure 6 (e)).
When $\mathcal{R}_{0}>1$, the prevalence of symptomatic and non-symptomatic infection increases with a more infectious disease (larger $\mathcal{R}_{0}^{D}$, Figure 6, (a), (b), (d)), although this increase is less noticeable when testing behaviour is strongly influenced by social interactions (larger $\mathcal{R}_{0}^{B}$). However, the prevalence of reported infections does not necessarily increase with increased disease contagiousness (Figure 6 (c)). Indeed, when testing behaviour is negligibly influenced by social interactions, we observe near zero reported infections in the steady state, providing a message of false-elimination when in fact infection rates are at their highest steady state value. When social contagiousness is increased for a fixed pathogen ($\mathcal{R}_{0}^{D}$), we observe an increase in reported cases followed by a decrease until elimination is reached ($\mathcal{R}_{0}<1$). This non-linear hill structure in reported cases of symptomatic infections indicates that it is important to know which side of the hill a population is on, in order to understand how to interpret reported case numbers correctly. For example, when testing rates are high and we observe a small proportion of the population testing positive for the infection, there is more confidence that the true case numbers throughout the population are reduced as well. However, if both testing rates and case numbers are low, there is a possibility that the actual burden of the infection on the population is much larger than it appears.
Figure 6: The effect of the social reproduction number and the behaviour-free reproduction number on endemic states. In each panel we display outputs for different combinations of the social reproduction number ($\mathcal{R}_{0}^{B}$, $y$-axis) and the behaviour-free reproduction number ($\mathcal{R}_{0}^{D}$, $x$-axis). The measures displayed are: (a) Total endemic infection state in the population ($O+A$); (b) Symptomatic endemic infection state in the population ($O$); (c) Observed symptomatic endemic infection state in the population ($T$); (d) Undetected symptomatic infection endemic state in the population ($I$); (e) Proportion of the population willing to test and isolate if symptomatic in the steady state of the population ($B$); (f) The basic reproduction number ($\mathcal{R}_{0}$). For (a)-(e), colour bars show the percentage of the population in each state. For (f), the colour bar shows the change in reproduction number. The dashed grey line shows $\mathcal{R}_{0}=1$ and the black cross shows the baseline parameters in Table 1.
4.2.3 Stability of steady states
There are four feasible steady state combinations for this BaD model for testing and isolation: no behaviour or disease (denoted $E_{00}$); no behaviour, disease endemic (denoted $E_{0D}$); behaviour endemic, no disease (denoted $E_{B0}$), and; behaviour and disease endemic (denoted $E_{BD}$). The stable regions of the steady states agree with the regions previously identified in [17] when $\omega_{1},\omega_{2},\omega_{3}>0$ (Figure 7 (a)) and when $\omega_{1}>0,\omega_{2}=\omega_{3}=0$ (Figure 7 (c)). In contrast, when $\omega_{3}=0$ and $\omega_{1},\omega_{2}>0$ we observe a non-linear boundary where the stable steady state switches from $E_{0D}$ to $E_{BD}$ (Figure 7 (b)). This boundary is driven by a non-linear relationship between social spread of behaviour ($\mathcal{R}_{0}^{B}$) and the perception of illness threat ($\omega_{2}$). This suggests that there are parameter combinations where an infection can be endemic in a population without being detectable in the sense that negligibly few tests are being conducted. Since those who test positive ($T$) are explicitly a subset of those willing to test and isolate ($B$), the threshold observed in Figure 7 (b) shows there is a need for enough people actively to seek testing and isolate if positive before the feedback mechanism of the perception of illness threat can take effect.
(a) $\omega_{1},\omega_{2},\omega_{3}>0$
(b) $\omega_{3}=0,\omega_{1},\omega_{2}>0$
(c) $\omega_{2}=\omega_{3}=0,\omega_{1}>0$
Figure 7: Stable regions of the steady states for different combinations of behavioural parameters. The $x$-axis shows the behaviour-free reproduction number ($\mathcal{R}_{0}^{D}$) varying through changes in $\beta$ and the $y$-axis shows the social reproduction number ($\mathcal{R}_{0}^{B}$) varying through changes in $\omega_{1}$. (a) Feasible steady states when all behavioural parameters are greater than zero ($\omega_{1},\omega_{2},\omega_{3}>0$). (b) Feasible steady states when $\omega_{1},\omega_{2}>0$ and $\omega_{3}=0$. (c) Feasible steady states when $\omega_{1}>0$ $\omega_{2}=\omega_{3}=0$. In each figure: dark blue shows the behaviour and disease free state $E_{00}$; light blue shows the no behaviour, disease endemic state $E_{0D}$; light red shows the behaviour endemic, no disease state $E_{B0}$, and; dark red shows the behaviour and disease endemic state $E_{BD}$.
5 Discussion
5.1 Findings in context
We have extended an existing behaviour and disease (BaD) model [17] to investigate the epidemiological consequences of the behavioural dynamics of symptomatic testing and isolation. Using our model, through theoretical derivations we have computed the basic reproduction number and identified stability conditions for epidemiological-behavioural steady states. Numerical simulation of epidemic and endemic regimes have then shown a variety of relationships between an epidemiological outcome of interest and the testing associated behavioural parameter values.
We have exemplified how BaD models enable us to explore the effects of interventions for reducing disease spread that are more targeted towards human behaviour. Previous individual-level transmission modelling studies have shown tracing and isolation alone are typically insufficient to control an epidemic [27]. Our results agree with these existing findings from a behavioural perspective,
Ultimately, the application and calibration of BaD transmission models to real-world problems is contingent on the availability of appropriate data sources [5, 28]. Compliance with testing policy is an important data input. COVID-19 testing interventions have shown that the testing behaviours, in multiple settings, have demographic and socioeconomic heterogeneities. In 2020, an analysis of the demographic determinants of testing incidence and COVID-19 infections in New York City neighbourhoods reported that people residing in poor or immigrant neighbourhoods were less likely to be tested, but the likelihood that a test was positive was larger in those neighbourhoods [29]. In Chile, test positivity and testing delays were higher in lower–socioeconomic status municipalities [30]. Mass testing data has also been used in conjunction with surveillance surveys to identify gaps in the uptake of public health interventions at both fine-scale levels and across sociodemographic groups in England [31] and Australia [32]. Improving our ability to onboard such data into BaD models, including real-time compliance estimates from community surveys [21], can result in more impactful testing interventions for disease control.
5.2 Study limitations and directions for further work
We regard the model we have presented as an entry point for quantitative study of the feedbacks between testing behaviour and infection dynamics. As with any model we acknowledge that it is, by necessity, a simplified representation of reality. It is important that we consider the modelling assumptions made and their potential limitations. Here we expand on the implications of our behavioural parameter assumptions, testing assumptions and reduction of model dimensionality via assuming the pathogen to have a single strain and there being a sole testing intervention deployed.
A core aspect of our presented model is parsimony. We therefore had fixed values for the behaviour effect parameters in the testing behaviour uptake rate ($\omega_{1},\omega_{2},\omega_{3}$) and testing behaviour abandonment rate ($\alpha_{1},\alpha_{2}$). Relaxing this assumption would permit temporal variation in these behavioural factors. One scenario of interest would be to have a decay rate on the perception of illness threat ($\omega_{2}$), corresponding to growing normality through time amongst the population for cases occurring and infection circulating in the community. Alternatively, we could vary the epidemiological metric to which the perception of illness threat is tied. We associated $\omega_{2}$ with those who tested positive. However, as there may be delays in reporting the count of positive tests to the wider population, different lags to this measure could be applied. Further, the case severity that the population responds to may alter during an outbreak (i.e. at different times the number of cases, hospitalisations and/or deaths may be attributed different importance); consideration of these behavioural feedbacks would require refinements to our model to include severe case outcomes (such as hospitalisations and deaths).
Another parsimony based assumption was that symptomatic ‘behaviour aware’ individuals who tested negative (due to test sensitivity being below 100%) had no change in their behaviour and were assumed to have the same infectiousness as those who were symptomatic and did not test. Removing this assumption would lead to more parameters being added to the model, increasing the parameter space to be explored beyond what we deem to be digestible in this single study. An ultimate ambition would be such processes being data-driven, meaning future analysis would benefit greatly from further behavioural data collection.
With regards to our testing assumptions, we implemented testing with idealised conditions of no delays in test accessibility and availability, or constraints on overall supply. In other words, we assumed that all those who wanted to test for infection could do so. Under circumstances where there were not sufficient tests both accessible and available, that would result in missed opportunities to identify infected individuals. If these individuals do not self-isolate during the infection episode (instead mixing with others in the population), we would anticipate higher peaks in infectious prevalence. These additional cases could then cascade into further onward transmission and new cases, putting further pressure on strained testing resources. Conceivably, issues with test accessibility and availability can have negative consequences on continued engagement with testing protocols for future infection episodes [33, 34]. These potential feedbacks between testing uptake and testing availability are an additional behavioural consideration that would benefit from data collection.
Lastly, we reduced the complexity of the modelled problem by treating the pathogen as only having a single strain and only considering one type of disease control (testing). The consideration of multiple strains/variants could be used to investigate the relationships to community survey data [21]. Furthermore, public health policies are often a combination of pharmaceutical (when available) and non-pharmaceutical interventions (NPIs). Multiple NPIs may also be used in conjunction with one another, such as social distancing and mask-usage in various geographic and social contexts during the COVID-19 pandemic [35, 36]. Due attention should be given to the parameterisation of the behaviour and epidemiological feedbacks in the presence of a package of interventions – the behavioural responses to testing may be influenced by the availability of other interventions – which is presently an open question and merits further research.
5.3 Conclusion
Models of infectious disease dynamics have traditionally been bereft of data-driven and/or theoretical knowledge of outbreak behavioural dynamics. Our study has provided a structured modelling approach for the joint consideration of mechanistic behaviour and disease transmission processes to the dynamics of (symptomatic) testing and isolation. To choose the most appropriate behavioural model and structure (with reasonable assumptions on the components of the behavioural model, to fit the target behaviour, infection and questions of interest), we encourage interdisciplinary collaboration across the behavioural, biological, data and mathematical sciences. With these data and modelling advancements we would like to build public trust in modelling studies, which can have onward benefits for the public health, economic and social response to future infectious disease outbreaks.
Acknowledgements
The authors would like to thank participants in the Isaac Newton Institute for Mathematical Sciences programme “Modelling and inference for pandemic preparedness” for helpful discussions on this work. TH would like to thank Istvan Kiss for sharing observations on the history of behaviour and disease modelling.
Financial disclosure
EMH is funded by The Pandemic Institute, formed of seven founding partners: The University of Liverpool, Liverpool School of Tropical Medicine, Liverpool John Moores University, Liverpool City Council, Liverpool City Region Combined Authority, Liverpool University Hospital Foundation Trust, and Knowledge Quarter Liverpool (EMH is based at The University of Liverpool). The views expressed are those of the author(s) and not necessarily those of The Pandemic Institute. TH is supported by the Wellcome Trust (Grant number 227438/Z/23/Z). MGR is supported by the Marsden Fund under contract 20-MAU-50. The authors would like to thank the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the programme “Modelling and inference for pandemic preparedness”, where work on this paper was initiated. This work was supported by EPSRC grant EP/Z000580/1. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
Conflicts of interest
All authors declare they have no conflict of interest.